Ben originally filled out 6 more applications than Henry. Then each boy filled out 4 additional applications, bringing the total to 28. How many applications did each boy originally fill out?




Answer by addingup(3677)

B = M+6
B+4+H+4 = 28
H+6+4+H+4 = 28
2H+14 = 28
2H = 14
H = 7
B = 7+6 = 13
Then each boy filled out 4 additional applications:
H = 7+4 = 11
B = 13+4 = 17
bringing the total to 28:
11+17 = 28 Correct